The income statement for the month of June, 2014 of Happy Smiles Enterprises contains the following information:

Revenues 7,000

Expenses:

Salaries and Wages Expense 3,000

Rent Expense 1,500

Advertising Expense 800

Supplies Expense 300

Insurance Expense 100

Total expenses 7,500

Net income 1,3000

The entry to close Income Summary to Owner’s, Capital includes

a. a debit to Revenues for $7,000.

b. credits to Expenses totalling $5,700.

c. a credit to Income Summary for $1,300

d. a credit to Owner's Capital for $1,300.
Answer #1

Answer : D) a Credit to Owers's Capital for $ 1,300

Net income = Sales Revenue- Total Expenses = 7000- ( 300+1500+800+300+100) = 7000-5700 = $1,300

Income Summary = $1,300

Close of income summary

Debit Credit
Income summary $ 1,300
Owner's Capital $1,300